# Improving Retrieval-Augmented Generation through Multi-Agent Reinforcement Learning

### v1

@@ 4.2. Detailed Configuration for Each Agent
4.2.1. Elements of Query Rewriter
Observation of Query Rewriter is defined as Equation (1), which contains prompt of Query Rewriter $Prompt_{QR}$ and the initial question $q$.
(1) $O_{QR}=\left\{Prompt_{QR},q\right\}$
Action Space of Query Rewriter corresponds to the vocabulary of LLMs $\mathcal{V}$.
(2) $A_{QR}=\mathcal{V}$
(3) $R_{QR}=R_{\text{shared}}+P_{QR}$
4.2.2. Elements of Selector
Observation of Selector is defined as Equation (4), which contains prompt of Selector $Prompt_{S}$, the initial question $q$ and the candidate documents set $D$ with $K$ documents.
(4) $O_{S}=\left\{Prompt_{S},q,D\right\}$
(5) $A_{S}=\left\{\text{"{0}"},\text{"{1}"},\ldots,\text{"{K-1}"},\text{"{Document}"},\text{"{,}"}\right\}$ ” }
(6) $R_{S}=R_{\text{shared}}+P_{S}$
4.2.3. Elements of Generator
Observation of Generator is in Equation (7), which contains prompt of Generator $Prompt_{G}$, the initial question $q$ and the selected candidate documents set $D_{\text{selected}}$ given by Selector.
(7) $O_{G}=\left\{Prompt_{G},q,D_{\text{selected}}\right\}$
Action Space of Generator $A_{G}$ is the same as Query Rewriter.
(8) $A_{G}=A_{QR}=\mathcal{V}$
Reward Function of Generator contains $R_{\text{shared}}$ and penalty term $P_{G}$, which is used to constrain the model from generating excessively long content. When the generated answer exceeds a certain length, $P_{G}$ is set to -0.5; otherwise, it is set to 0.
(9) $R_{G}=R_{\text{shared}}+P_{G}$
4.3. Warm Start with SFT
In preparation for joint optimization of multiple modules using Multi-Agent PPO, it is essential to perform a warm start for each trainable module. The warm start enables the model to better adhere to instructions across diverse tasks and reduces the exploration space during MARL joint training, thereby enhancing the efficiency of exploration and exploitation. Within the MMOA-RAG framework, there are three trainable modules: the Query Rewriter, the Selector, and the Generator. Consequently, it is necessary to construct training data for the SFT of each corresponding task.
4.3.1. Query Rewriter
@@
In Rewrite-Retrieve-Read (Ma et al., 2023), a small language model was trained using PPO to effectively rewrite queries for RAG. Building on this approach, we utilize the publicly available query rewriting data from Rewrite-Retrieve-Read as the SFT dataset to warm start the Query Rewriter in MMOA-RAG.
4.3.2. Selector
@@
The task of the Selector is to choose a subset $D_{\text{selected}}$ that are helpful for answering a question from a given set $D$ with $K$ candidate documents. The output format of the Selector is the IDs of the documents in $D_{\text{selected}}$ (e.g., Document0, Document4, Document6, Document7), as shown in the prompt of Selector in Table 1. Therefore, to construct SFT data for the Selector, the ground truth should be the IDs of documents that are truly useful for answering the question. One method to obtain the ground truth is to employ advanced LLMs, such as GPT-4o, to provide the ground truth. However, we have found that this approach does not yield results as good as expected. Additionally, BGM (Ke et al., 2024) introduced and optimized a bridge module which is similar to the Selector module. They proposed a method called synthesis silver passage sequence (Synthesis SPS) to construct the ground truth for SFT data. However, the Synthesis SPS method requires examining each candidate document $d_{i,j}\in D_{i}$ (candidate documents of question $i$), invoking the LLM for each check, and comparing the utility values before and after the check, making it a complex and costly method.
Table 1. The prompt of Selector agent.
@@
Figure 2. The convenient approach to construct the SFT data for Selector.
4.3.3. Generator
The Generator is responsible for producing the final answer, $Ans_{\text{predict}}$, based on the $D_{\text{selected}}$ provided by the Selector. Therefore, the ground truth for the SFT data of Generator is the golden answer $Ans_{\text{golden}}$.
@@ 4.3.3. Generator
(10) $\mathcal{L_{\text{SFT}}}(\theta)=-\sum_{n=1}^{N}\log P(Y_{i}\mid X_{i};\theta)$
@@
4.4. Multi-Agent Optimization
[…]
In the multi-agent optimization process, there are three models to consider: the Actor model, the Critic model, and the SFT model. The parameters for these models are denoted as $\theta$, $\phi$, and $\theta_{\text{SFT}}$, respectively. The role of the Actor model is to provide the response $Answer_{i}$ based on the observation $O_{i}$ for each agent $i$. The Critic model is responsible for estimating the state-value function $V_{\phi}^{i,t}$, which is a classic setup in Actor-Critic architecture within RL algorithms. The SFT model serves as a baseline for the Actor model, similar to InstructGPT (Ouyang et al., 2022). The objective is to update the parameters of both the Actor and Critic models. The overall loss function, $\mathcal{L}(\theta,\phi)$, consists of two terms: $\mathcal{L}_{\text{Actor}}(\theta)$ and $\mathcal{L}_{\text{Critic}}(\phi)$:
(11) $\mathcal{L}(\theta,\phi)=\mathcal{L}_{\text{Actor}}(\theta)+\alpha*\mathcal{L}_{\text{Critic}}(\phi)$
The Actor loss function presented in Equation (12) is similar to that used in the typical single-agent PPO (Schulman et al., 2017) algorithm. The primary difference is that multiple agents are being optimized. In Equation (12), $i\in\mathcal{G}$ denotes the three agents: Query Rewriter, Selector, and Generator. The term $r_{t}^{i}$ in Equation (13) denotes the importance sampling ratio, which measures the difference between the new and old policies. The expression $\hat{A}_{\pi_{\theta}}^{i,t}$ in Equation (14) is the advantage function, estimated using Generalized Advantage Estimation (GAE) (Schulman et al., 2015). The variable $\delta_{t}^{i}$ in Equation (15) is known as the temporal difference (TD) error at time step $t$.
(12) $\mathcal{L_{\text{Actor}}}(\theta)=\sum_{i}\sum_{t}\min\left(r^{i}_{t}\hat{A}_{{\pi}_{\theta}}^{i,t},\ \text{clip}\left(r^{i}_{t},1-\epsilon,1+\epsilon\right)\hat{A}_{{\pi}_{\theta}}^{i,t}\right)$
(13) $r^{i}_{t}=\frac{\pi_{\theta}(a^{i}_{t}\mid s^{i}_{t})}{\pi_{\theta_{\text{old}}}(a^{i}_{t}\mid s^{i}_{t})}$
(14) $\hat{A}_{{\pi}_{\theta}}^{i,t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{i}_{t+l}$
(15) $\delta^{i}_{t}=R(s^{i}_{t},a^{i}_{t})+\gamma V_{\phi}(s^{i}_{t+1})-V_{\phi}(s^{i}_{t})$
Similar to InstructGPT (Ouyang et al., 2022), the final reward function $R(s^{i}_{t},a^{i}_{t})$ is defined in Equation (16). The distinction is that our approach does not require a trained reward model, as we use the evaluation metric (F1 score) of the predicted answers $Ans_{\text{predict}}$ of Generator as the shared reward $R_{\text{shared}}$ for all agents. The penalty term $P_{i}$ can also be easily obtained from the output of each agent, as introduced in Section 4.2. The components $R_{i}$ in Equation (16) are defined in Equations (3), (6), or (9). And $Answer_{i}$ represents the output generated by each agent $i$ based on its individual observation $O_{i}$.
[…]
(17) $\mathcal{L}_{\text{Critic}}(\phi)=\sum_{i}\sum_{t}\max\left[(\Delta V_{i,t})^{2},\left(\text{clip}\left(V_{\phi}^{i,t},V_{\phi_{\text{old}}}^{i,t}\pm\epsilon\right)-V_{\text{target}}^{i,t}\right)^{2}\right]$

### v2

@@ 3.4.2 Multi-Agent Optimization
In the multi-agent optimization process, there are three models to consider: the Actor model, the Critic model, and the SFT model. The parameters for these models are denoted as $\theta$, $\phi$, and $\theta_{\text{SFT}}$, respectively. The role of the Actor model is to provide the response $Answer_{i}$ based on the observation $O_{i}$ for each agent $i$. The Critic model is responsible for estimating the state-value function $V_{\phi}^{i,t}$, which is a classic setup in Actor-Critic architecture within RL algorithms. The SFT model serves as a baseline for the Actor model, similar to InstructGPT ouyang2022training . The objective is to update the parameters of both the Actor and Critic models. The overall loss function, $\mathcal{L}(\theta,\phi)$, consists of two terms: $\mathcal{L}_{\text{Actor}}(\theta)$ and $\mathcal{L}_{\text{Critic}}(\phi)$:
$\mathcal{L}(\theta,\phi)=\mathcal{L}_{\text{Actor}}(\theta)+\alpha*\mathcal{L}_{\text{Critic}}(\phi)$ (10)
The Actor loss function presented in Equation (11) is similar to that used in the typical single-agent PPO schulman2017proximal algorithm. The primary difference is that multiple agents are being optimized. In Equation (11), $i\in\mathcal{G}$ denotes the three agents: Query Rewriter, Selector, and Generator. The term $r_{t}^{i}$ in Equation (12) denotes the importance sampling ratio, which measures the difference between the new and old policies. The expression $\hat{A}_{\pi_{\theta}}^{i,t}$ in Equation (13) is the advantage function, estimated using Generalized Advantage Estimation (GAE) schulman2015high . The variable $\delta_{t}^{i}$ in Equation (14) is known as the temporal difference (TD) error at time step $t$.
$\mathcal{L_{\text{Actor}}}(\theta)=\sum_{i}\sum_{t}\min\left(r^{i}_{t}\hat{A}_{{\pi}_{\theta}}^{i,t},\ \text{clip}\left(r^{i}_{t},1-\epsilon,1+\epsilon\right)\hat{A}_{{\pi}_{\theta}}^{i,t}\right)$ (11)
$r^{i}_{t}=\frac{\pi_{\theta}(a^{i}_{t}\mid s^{i}_{t})}{\pi_{\theta_{\text{old}}}(a^{i}_{t}\mid s^{i}_{t})}$ (12)
$\hat{A}_{{\pi}_{\theta}}^{i,t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{i}_{t+l}$ (13)
$\delta^{i}_{t}=R(s^{i}_{t},a^{i}_{t})+\gamma V_{\phi}(s^{i}_{t+1})-V_{\phi}(s^{i}_{t})$ (14)
Similar to InstructGPT ouyang2022training , the final reward function $R(s^{i}_{t},a^{i}_{t})$ is defined in Equation (15). The distinction is that our approach does not require a trained reward model, as we use the evaluation metric (F1 score) of the predicted answers $Ans_{\text{predict}}$ of Generator as the shared reward $R_{\text{shared}}$ for all agents. The penalty term $P_{i}$ can also be easily obtained from the output of each agent, as introduced in Section 3.3. The components $R_{i}$ in Equation (15) are defined in Equations (3), (6), or (9). And $Answer_{i}$ represents the output generated by each agent $i$ based on its individual observation $O_{i}$.
$R(s^{i}_{t},a^{i}_{t})=\begin{cases}0,&\text{if }t<T\\ \underbrace{R_{\text{shared}}+P_{i}}_{R_{i},\text{ and }i\in\mathcal{G}}-\beta\log\left(\frac{\pi_{\theta}(\mathit{Answer}_{i}\mid O_{i})}{\pi_{\theta_{\text{SFT}}}(\mathit{Answer}_{i}\mid O_{i})}\right),&\text{if }t=T\end{cases}$ (15)
$\mathcal{L}_{\text{Critic}}(\phi)=\sum_{i}\sum_{t}\max\left[(\Delta V_{i,t})^{2},\left(\text{clip}\left(V_{\phi}^{i,t},V_{\phi_{\text{old}}}^{i,t}\pm\epsilon\right)-V_{\text{target}}^{i,t}\right)^{2}\right]$ (16)
The pseudocode for multi-agent optimization based on MAPPO is shown in Algorithm 1 in Appendix C, which corresponds to the overall framework of MMOA-RAG depicted in Figure 1. For a specific question, the first step is to execute the Collect Rollout process. This process involves passing through the Query Rewriter, Retriever, Selector, and Generator, and the computed tuple $\mathcal{T}=\left((O_{QR},subq,R_{QR}),(O_{S},IDs,R_{S}),(O_{G},Ans_{\text{predict}},R_{G})\right)$ is stored in the replay buffer $\mathcal{M}$. Next, the Policy and Value Optimization process is executed where the GAE is used to estimate the advantage function $\hat{A}_{{\pi}_{\theta}}^{i,t}$. Subsequently, the overall loss function $\mathcal{L}(\theta,\phi)$ is calculated, and the parameters of both the Actor and Critic models are updated. Additionally, to accelerate the entire training process, we can run a minibatch in parallel. Ultimately, we obtain a well-trained Actor model used for subsequent inference and evaluation.
4 Experiments
RQ1: How does our MMOA-RAG perform compared to existing RAG optimization methods?
RQ2: How does the joint optimization of individual modules in the RAG pipeline contribute to the effectiveness of MMOA-RAG framework?
RQ3: Can MMOA-RAG exhibit generalizability across different RAG systems?
4.1 Experimental Settings
Datasets and Evaluation We conducted experiments using MMOA-RAG alongside various baseline models across three open-domain QA datasets: HotpotQA yang2018hotpotqa , 2WikiMultihopQA ho2020constructing , and AmbigQA min2020ambigqa . The candidate documents are all retrieved from Wikipedia passages for three datasets. We employ three key evaluation metrics—Accuracy, Exact Match (EM), and F1 score—to assess the performance of the RAG methods.
Implementation Details We utilize Contriever izacard2021unsupervised as the Retriever for most experiments. And the Selector consistently receives a fixed set of $K=10$ documents as input. Besides, we employ Llama-3-8B-Instruct dubey2024llama as the foundational LLM for all the baselines and MMOA-RAG.
We compare MMOA-RAG with different baseline methods: LLM w/o RAG, Vanilla RAG w/o train, Vanilla RAG w SFT, SELF-RAG asai2023self , RetRobust yoran2023making , Rewrite-Retrieve-Read ma2023query , BGM ke2024bridging , RAG-DDR li2024rag .
The detailed introduction of experimental settings and baselines can be seen in Appendix D.
4.2 Comparisons with Other Methods
Table 1: Performance for different methods across datasets. All the results in this table are obtained using Contriever izacard2021unsupervised as the retrieval model. In each dataset, the highest baseline value is underscored. The symbol $\Delta$ displays the improvement of MMOA-RAG over the best baseline.
Methods HotpotQA 2WikiMultihopQA AmbigQA
Acc EM F1 Acc EM F1 Acc EM F1
LLM w/o RAG 25.08 21.31 31.18 27.78 23.68 29.47 27.21 20.96 33.42
Vanilla RAG w SFT 36.18 32.30 44.49 39.47 38.28 43.36 34.41 30.74 44.36
SELF-RAG asai2023self 30.42 27.77 38.93 36.32 35.39 38.86 28.35 25.70 39.04
RetRobust yoran2023making 37.69 34.60 46.49 41.02 39.73 44.51 35.13 32.37 44.78
Rewrite-Retrieve-Read ma2023query 38.03 33.93 46.32 40.40 39.17 44.17 35.94 31.90 45.92
BGM ke2024bridging 36.05 32.76 44.54 39.61 38.61 43.29 36.01 32.53 45.76
RAG-DDR li2024rag 35.20 32.65 44.26 40.49 39.45 44.18 36.25 32.55 45.83
MMOA-RAG (ours) 39.15 36.15 48.29 42.73 41.52 46.40 38.85 34.75 48.59
$\Delta$ +1.12 +1.55 +1.80 +1.71 +1.79 +1.89 +2.60 +2.20 +2.67
We conducted a comparative analysis of MMOA-RAG against multiple baselines, with the results presented in Table 1. To ensure fairness in comparison, all methods utilized Llama-3-8B-Instruct as the backbone LLM, and all baselines were re-implemented according to the settings delineated in Appendix D.2.
In addition, we tested various methods based on other retriever, BGE xiao2024c and E5 wang2022text , and the results can be seen in Table 7 in Appendix E. And we also perform out-of-domain experiments, the results can be seen in Table 8 in Appendix F. In summary, The results in Table 1, Table 7, Table 8, and the analysis in this Section 4.2 jointly answer the RQ1. And the case study can be found in Appendix I, from which we can intuitively understand the advantages of multi-module joint optimization.
4.3 Ablation Experiments on the Optimization of Different Agents
To demonstrate the necessity of multi-agent joint optimization in RAG systems, we present ablation experiments in this section. The MMOA-RAG framework, depicted in Figure 1, consists of three agents: $i\in\{\text{Query Rewriter (QR), Selector (S), Generator (G)}\}$. In Figure 2, MMOA-RAG w/o $i$ denotes the variant where agent $i$ is excluded from the complete optimization process of multi-agent joint optimization.
(a) HotpotQA
(b) 2WikiMultihopQA
(c) AmbigQA
Figure 2: Ablation about Optimizing Different Agents. In this figure, MMOA-RAG w/o $i$ ($i\in\{\text{QR, S, G}\}$) denote the variant where agent $i$ is excluded from the complete optimization process of multi-agent joint optimization.
Figure 3: Ablation experiments on AmbigQA dataset. The horizontal axis represents the number of training samples, while the vertical axis denotes the shared reward $R_{\text{shared}}$ (F1 score) during the training process.
As illustrated in Figure 2, the complete version of MMOA-RAG, where all three modules are jointly optimized, delivers the highest performance. This underscores the effectiveness of multi-agent joint optimization within the RAG system and validates the importance of optimizing multiple modules concurrently. Additionally, the MMOA-RAG w/o S variant achieves the best performance among the three ablation configurations. The Selector’s primary function is to refine the candidate document set $D$, yielding a higher-quality subset $D_{\text{selected}}$, which enhances the Generator’s ability to produce a superior answer $Ans_{\text{predict}}$. However, through the joint optimization by MAPPO, the Generator acquires some denoising capabilities. Consequently, satisfactory results can be achieved even when the Selector is not optimized during joint optimization.
We also present the trajectory of the shared reward $R_{\text{shared}}$ during the training process based on ablation experiments conducted on the AmbigQA dataset, as shown in Figure 3. From Figure 3, it is evident that the reward curve for MMOA-RAG demonstrates the fastest convergence rate and achieves the highest final convergence value. This underscores the effectiveness of joint optimization across multiple modules in significantly and efficiently enhancing the performance of the RAG system. Furthermore, the training curve for MMOA-RAG w/o G in Figure 3 is noticeably slower compared to other algorithms, and the test results for MMOA-RAG w/o G on the AmbigQA dataset, as shown in Figure 2, are the poorest. These findings suggest that the Generator module is the most critical component for the single-hop AmbigQA dataset.
The results in Figure 2 and Figure 3 answer the RQ.2 that it is more effective to optimize multiple modules in a RAG system simultaneously.
4.4 Generality Experiments on RAG Systems with Varying Module Configurations
Table 2: Generality Experiments on RAG Systems with Varying Module Configurations. In the second column, SFT and MAPPO refer to the current module configuration following the warm start training stage (Section 3.4.1) and the MAPPO joint training stage (Section 3.4.2), respectively. The symbol $\Delta$ signifies the enhancement achieved in the MAPPO stage relative to the SFT stage.
Modules Training Stage & Delta HotpotQA 2WikiMultihopQA AmbigQA
Acc EM F1 Acc EM F1 Acc EM F1
QR+S+G SFT 36.00 33.04 44.69 39.54 38.50 42.97 36.55 32.60 46.71
MAPPO 39.15 36.15 48.29 42.73 41.52 46.40 38.85 34.75 48.59
$\Delta$ +3.15 +3.11 +3.60 +3.19 +3.02 +3.43 +2.30 +2.15 +1.88
S+G SFT 34.25 32.18 43.14 38.93 37.97 42.40 35.85 32.35 45.82
MAPPO 38.23 34.85 47.07 41.79 40.57 45.25 37.60 33.90 47.19
$\Delta$ +3.98 +2.67 +3.93 +2.86 +2.60 +2.85 +1.75 +1.55 +1.37
QR+G SFT 36.76 32.78 45.00 39.15 37.89 42.91 35.50 31.50 45.31
MAPPO 38.90 35.89 47.94 42.43 41.01 46.19 37.65 33.50 47.53
$\Delta$ +2.14 +3.11 +2.94 +3.28 +3.12 +3.28 +2.15 +2.00 +2.22
5 Conclusions and Future Works
In future works, we intend to explore the application of MMOA-RAG in more complex workflows. This exploration will include scenarios where the RAG workflows are organized as a directed acyclic graph, as well as situations involving dynamic workflows within agentic RAG. Additionally, it is important to assess the cost and latency associated with specific modules within the RAG system. In this regard, the design of the reward function should not be exclusively based on evaluation metrics, such as the F1 score highlighted in this paper. Instead, it should aim to strike a balance between effectiveness and cost in those partially cooperative RAG scenarios.
Appendix
Appendix A Detailed Related Works
A.1 End-to-end Optimization in OpenQA
ORQA [24] is an open-domain QA system that learns end-to-end evidence retrieval and answer generation using only question-answer pairs, enabled by pretraining with an Inverse Cloze Task. REALM [12] is an end-to-end optimizing framework that enhances language model pre-training with a retrieval-augmented approach. [25] introduces Retrieval-Augmented Generation as RAG, a model that combines pre-trained language models with non-parametric memory for improved performance on knowledge-intensive NLP tasks. [17] propose a knowledge distillation method to train retriever models using synthetic labels derived from reader model attention scores. Stochastic RAG [54] introduces a novel end-to-end optimization framework for RAG through expected utility maximization.
A.2 RAG without Parameters Updating
These methods typically involve designing a novel RAG mechanism to enhance the performance of LLMs on question answering tasks. For example, DSP [23] leverages sophisticated interactions between retrieval and language models to address knowledge-intensive NLP tasks. FLARE [20], an active retrieval augmented generation method, enhances text generation by dynamically retrieving relevant information throughout the process, showing superior performance across various long-form knowledge-intensive tasks. ITER-RETGEN [39] is an iterative retrieval-generation synergy method that enhances retrieval-augmented large language models by synergistically combining retrieval and generation in an iterative manner. Search-in-the-Chain [49] is a framework that interactively enhances Large Language Models with search capabilities to improve performance on complex, knowledge-intensive tasks. SELF-RAG [1] enhances language model quality and factuality through self-reflective retrieval and generation. DRAGIN [43] is a dynamic RAG framework that addresses the real-time information needs of LLMs during text generation, enhancing performance on knowledge-intensive tasks. GenGround [40] synergizes large language model knowledge with external documents to enhance multi-hop question answering through an iterative process of generating answers and grounding them in evidence. Astute RAG [45] is a approach that enhances the robustness of Retrieval-Augmented Generation for Large Language Models by adaptively integrating internal and external knowledge while resolving knowledge conflicts.
A.3 RAG with Parameters Updating
A.3.1 Optimizing RAG with SFT
INFO-RAG [50], an unsupervised training method, enhances the capacity of large language models to integrate and refine information from retrieved texts. LongRAG [55] introduces a dual-perspective retrieval-augmented generation system to enhance understanding of complex long-context knowledge for improved performance in long-context question answering tasks. In INSTRUCTRAG [47], generation accuracy and trustworthiness are enhanced by explicitly denoising retrieved information through self-synthesized rationales, outperforming standard RAG approaches without additional supervision.
A.3.2 Optimizing RAG with RL
Some existing works use PPO [38] algorithm to fine-tune LLMs. In Rewrite-Retrieve-Read framework [29], a small language model is trained with reinforcement learning to rewrite queries for RAG. BGM [22] proposes a novel bridge mechanism between retrieval model and LLMs and uses PPO to optimize the parameters of the bridge to filter for more helpful documents. SMARTRAG [9] optimizes an iterative RAG framework with reward, which includes a decision maker and a policy network. RAG-Star [19] is a reasoning approach that combines Monte Carlo Tree Search (MCTS) to improve the complex reasoning abilities of LLMs. Additionally, concurrent work Search-r1 [21] and R1-Searcher [42] both use answer-based reward to improve the reasoning in RAG.
Some other works use DPO [32] or similar alignment algorithms to optimize LLMs. A noise-filtering method [60] is proposed by optimizing mutual information between compressed data and output while minimizing it with the retrieved passage. ATM [59], an Adversarial Tuning Multi-agent system, enhances the robustness and performance of retrieval-augmented generators in question answering by iteratively tuning against an adversarial attacker agent to better discriminate useful documents and resist fabricated content. SEER [57] proposes a novel self-aligned evidence extraction learning framework aimed at enhancing RAG performance by optimizing the extraction of high-quality, concise, and relevant evidence. RAG-DDR [27] optimizes RAG systems by aligning data preferences between modules through DDR, resulting in enhanced performance on knowledge-intensive tasks. MAO-ARAG [5] proposes a hierarchical RAG framework that achieves a balance between effectiveness and efficiency in RAG by dynamically orchestrating executors through optimizing the planner agent with RL, thereby achieving Pareto optimality.
Appendix B How to Construct Training Data for the SFT Training Process
Before the multi-module joint training process of MAPPO, the warm-up to each trainable modules is necessary. The warm-up can make LLMs to follow the instructions better and reduce the exploration space for the MAPPO training process to stabilize and accelerate the joint training process.
In this section, we introduce the details of constructing training data for each modules.
B.1 Query Rewriter
Query Rewriter is to reformulate the initial query $q$, which may be too complex or ambiguous to resolve with a single retrieval, into a set of sub-questions denoted as subqsubq. The prompt of Query Rewriter is in Table 3.
@@ B.1 Query Rewriter
In Rewrite-Retrieve-Read [29], a small language model was trained using PPO to effectively rewrite queries for RAG. Building on this approach, we utilize the publicly available query rewriting data from Rewrite-Retrieve-Read as the SFT dataset to warm start the Query Rewriter in MMOA-RAG.
Table 3: The prompt of Query Rewriter agent.
system: You are a professional assistant skilled at rewriting complex or unclear questions into simpler, more searchable subquestions. assistant: Okay, I will provide the rewritten sub-questions.
user: Please help me rewrite or decompose the given questions into sub-questions, making them easier to search for answers in a search engine. The rewritten sub-questions must have logical connections and dependencies, without being overly repetitive in meaning. Additionally, avoid using vague demonstrative pronouns and similar terms.
assistant: Okay, I will provide the rewritten sub-questions.
user: Original question is {content of Question}. Now rewrite or decompose the original question into sub-questions according to the above requirements, and only output the rewritten subquestions in the format of one subproblem per line without any additional content. Additionally, avoid using vague demonstrative pronouns and similar terms, avoid the duplicate subquestions.
B.2 Selector
The task of the Selector is to choose a subset $D_{\text{selected}}$ that are helpful for answering a question from a given set $D$ with $K$ candidate documents.
@@ B.2 Selector
The output format of the Selector is the IDs of the documents in $D_{\text{selected}}$ (e.g., Document0, Document4, Document6, Document7), as shown in the prompt of Selector in Table 4. Therefore, to construct SFT data for the Selector, the ground truth should be the IDs of documents that are truly useful for answering the question. One method to obtain the ground truth is to employ advanced LLMs, such as GPT-4o, to provide the ground truth. However, we have found that this approach does not yield results as good as expected. Additionally, BGM [22] introduced and optimized a bridge module which is similar to the Selector module. They proposed a method called synthesis silver passage sequence (Synthesis SPS) to construct the ground truth for SFT data. However, the Synthesis SPS method requires examining each candidate document $d_{i,j}\in D_{i}$ (candidate documents of question $i$), invoking the LLM for each check, and comparing the utility values before and after the check, making it a complex and costly method.
Table 4: The prompt of Selector agent.
@@
Document1: {content of Document1}
Document(K-2): {content of Document(K-2)}
@@
Figure 4: The convenient approach to construct the SFT data for Selector.
B.3 Generator
The Generator is responsible for producing the final answer, $Ans_{\text{predict}}$, based on the $D_{\text{selected}}$ provided by the Selector. Therefore, the ground truth for the SFT data of Generator is the golden answer $Ans_{\text{golden}}$. And the prompt of Generator is in Table 5.
Table 5: The prompt of Generator agent.
system: You are a helpful, respectful and honest assistant. Your task is to predict the answer to the question based on the given documents. If you don’t know the answer to a question, please don’t share false information. Answer the question as accurately as possible. assistant: Okay, I will provide the answer to the question based on the corresponding documents. Please provide the question and the corresponding documents.
user: Question: {content of Question}
Document0: {content of Document0}
Document3: {content of Document3}
Document7: {content of Document7}
assistant: OK, I received the Question and the corresponding Documents.
user: Given the Question and the corresponding Documents, predict the answer to the Question as briefly and accurately as possible based on the Documents. Only give the brief and accurate answer with the form of **answer** and nothing else.
@@ B.3 Generator
$\mathcal{L_{\text{SFT}}}(\theta)=-\sum_{n=1}^{N}\log P(Y_{i}\mid X_{i};\theta)$ (17)
@@
Appendix C The Pseudocode of Multi-Agent Training Process of MMOA-RAG
Algorithm 1 is the pseudocode for multi-agent optimization based on MAPPO.